# Samer Al-Hamadani

### v1

@@ 5.1 Novel Category Evaluation with Tiered Web Representation
Novel category evaluation reveals fundamental architectural differences with critical insights from stratified web prevalence sampling. YOLOv8m detects exactly 0 of 200 novel product instances across all three tiers as architecturally expected, validating that supervised detector cannot generalize beyond training taxonomies regardless of web representation density, model quality, or training data scale.
@@
In stark contrast, Gemini successfully detects 105 of 200 instances (52.3% overall accuracy, CI [45.4%, 59.1%]) with dramatic performance variation across web prevalence tiers:
@@
Tier 1 (Highly Web-Prevalent, N=100): 79.0% accuracy [70.1%, 86.2%] This tier includes consumer electronics with millions of online images (AirPods, Tesla, smart home devices). High performance reflects substantial VLM pre-training exposure through product marketing, user-generated content, e-commerce listings, and review platforms. Per-category breakdown:
@@
  • AirPods Pro: 90% (9/10) - Distinctive charging case, ubiquitous Apple marketing
@@
  • Ring Doorbell: 90% (9/10) - Camera module prominent, extensive installation videos
@@
  • Tesla Model 3: 80% (8/10) - Minimalist design, automotive media coverage
@@
  • Amazon Echo Dot: 80% (8/10) - Cylindrical shape, tech review prevalence
@@
  • Peloton Bike: 80% (8/10) - Screen mount distinctive, fitness media coverage
  • Google Nest Hub: 70% (7/10) - Display device similarities to tablets
  • Oculus Quest 2: 80% (8/10) - VR visor distinctive, gaming media coverage
  • DJI Mavic Drone: 70% (7/10) - Drone design similarities, tech enthusiast coverage
  • Nintendo Switch OLED: 80% (8/10) - Gaming console prevalent, product launches
  • Apple AirTag: 70% (7/10) - Small size challenging, tracking device niche
@@
Tier 2 (Moderately Web-Prevalent, N=50): 48.0% accuracy [34.2%, 62.1%] This tier includes products with thousands rather than millions of online images. Moderate performance reflects partial VLM exposure:
@@
  • Electric Scooters: 60% (6/10) - Vehicle-like, urban mobility coverage
  • N95 Masks: 50% (5/10) - Medical equipment similarity, pandemic surge imagery
  • Standing Desks: 50% (5/10) - Furniture similarities, office equipment niche
@@
  • Smart Door Locks: 40% (4/10) - Hardware similarities, home security niche
@@
  • Portable Power Stations: 40% (4/10) - Battery pack similarities, camping niche
@@
Tier 3 (Rare Specialized Equipment, N=50): 30.0% accuracy [18.1%, 43.7%] This tier includes equipment with fewer than 500 online images, testing genuine zero-shot capability. Low performance confirms limited pre-training exposure:
@@
  • Industrial 3D Metal Printers: 20% (2/10) - Specialized manufacturing, confidential
@@
  • Underwater ROV Equipment: 30% (3/10) - Marine research niche, limited documentation
@@
  • Precision Laboratory Pipettes: 40% (4/10) - Scientific equipment, lab photography rare
  • Professional Broadcast Cameras: 30% (3/10) - Specialized models, production environments
@@
  • Medical Ultrasound Transducers: 30% (3/10) - Clinical equipment, privacy restrictions
@@
These tiered results validate our hypothesis: VLM zero-shot performance directly correlates with web representation density, ranging from near-supervised levels (79%) on web-prevalent products to genuine zero-shot challenges (30%) on rare equipment. However, even 30% substantially exceeds YOLO’s architectural impossibility (0%), establishing decisive qualitative advantage for applications encountering any objects outside training distributions.
Figure 3 visualizes novel category performance illustrating the relationship between visual characteristics and zero-shot detection accuracy.
Figure 3: Novel category zero-shot detection performance.
@@ 6 Cost-Effectiveness Analysis and Deployment Scenarios
Integration of technical performance results with cost modeling framework enables comprehensive economic analysis revealing when each paradigm achieves optimal return on investment. Table 3 presents TCO projections and Cost per Correct Detection calculations at representative inference volumes for 100-category systems, demonstrating how cost-effectiveness evolves with deployment scale.
Table 3: Comparative inference cost and efficiency between YOLO and Gemini
Inference YOLO YOLO Gemini Gemini CCD Best
Volume TCO (\$) CCD (\$) TCO (\$) CCD (\$) Advantage Choice
@@
[…]
150,000,000 65,232 0.00048 37,500 0.00052 YOLO better YOLO
200,000,000 82,832 0.00046 50,000 0.00069 YOLO better YOLO
[…]
These break-even volumes far exceed typical deployment scales for most organizations. A startup processing 1,000 images daily reaches only 365,000 annual inferences, operating at 0.7% of TCO break-even and 0.3% of CCD break-even. A small-to-medium business processing 5,000 images daily reaches 1.825 million annually (3.3% of break-even). Even substantial enterprises processing 50,000 images daily require 3.0 years continuous operation to approach 55 million inference TCO break-even, with realistic deployments facing evolving requirements and hardware refreshes preventing indefinite amortization. The practical implication establishes that supervised training proves economically rational only for the highest-volume production systems processing hundreds of thousands to millions of images daily across stable category taxonomies.
@@
Scenario Daily Vol Cat-egories Budget Accu-racy Late-ncy Optimal Primary Rationale
@@
Startup E-commerce 1,000 50 <<\$5K 65%+ Any Gemini YOLO \$6,020 upfront exceeds budget; Gemini \$0 upfront + \$90/year enables immediate deployment
@@
SMB Retail Analytics 5,000 100 \$10K 75%+ <<1s Gemini Annual vol 1.8M far below 55M break-even; Gemini \$456/year vs YOLO \$11,616 upfront
@@
Research Wildlife 333 500+ <<\$3K 60%+ Any Gemini YOLO \$180K annotation impossible; novel species require zero-shot; Gemini \$30/year viable
@@
Medical Imaging 10,000 12 \$50K 85%+ <<10s Hybrid Expert annotation \$43K prohibitive; hybrid VLM screening + YOLO verification balances cost/accuracy
@@
Enterprise Inventory 500,000 200 \$500K 90%+ <<50ms YOLO Annual vol 182.5M exceeds break-even; YOLO 5-yr \$75K vs Gemini \$228K; accuracy+latency critical
@@
Autonomous Vehicles 10M 20 Unlim-ited 95%+ <<20ms YOLO Safety demands 95%+ accuracy; 9.2ms latency mandatory; 287ms Gemini architecturally unsuitable
@@
An e-commerce startup deploying detection across 50 product categories initially, with planned expansion to additional categories, processing approximately 1,000 images daily faces YOLO upfront costs of \$6,020 (annotation \$5,400 + training \$120 + hardware \$500) versus zero for Gemini. Note that most startups begin with 5–10 core categories and expand incrementally; our 50-category scenario represents a more aggressive initial deployment. With annual operational expenses of \$6,020 for supervised versus \$90 for zero-shot (365K images $\times$ \$0.00025 minus free-tier coverage), the supervised approach exceeds typical startup budgets under \$5,000 while requiring 3–4 weeks for annotation completion, making Gemini the economically viable option for rapid market entry. Gemini delivers 72.8% accuracy exceeding typical 65% threshold for catalog applications where human review provides backup verification. Additionally, category evolution at 5–10 new products monthly imposes \$144 per category (\$90 annotation + \$54 training) for YOLO versus \$0 for Gemini text queries, further strengthening zero-shot advantages for rapidly evolving taxonomies common in e-commerce.
@@
Small-to-medium businesses operating retail analytics across multiple stores with 5,000 images daily face annual volume of 1.825 million images, reaching only 3.3% of the 55.3 million break-even thresholds. YOLO requires \$11,616 upfront investment for 100-category system versus Gemini annual cost of \$456 (after free-tier deduction), yielding 25-fold cost advantage for zero-shot approaches. While YOLO provides superior 89.7% accuracy versus Gemini 72.8%, the 17 percentage point improvement costs \$11,160 additional over first year equivalent to \$657 per percentage point of accuracy gain. For retail analytics applications where 75% accuracy suffices with occasional manual verification, this premium proves economically unjustifiable.
Research laboratories exploring wildlife monitoring face annotation costs scaling with species count. For a realistic initial study targeting 50 species with 10,000 monthly images (333 daily average), annotation costs reach \$18,000 (50 categories $\times$ 100 images $\times$ 3 boxes $\times$ \$0.30 $\times$ 1.20), already exceeding typical grant budgets of \$5,000–10,000 for pilot projects. Expanding to comprehensive biodiversity monitoring across 500+ species would require \$180,000 annotation investment, prohibitive for most research contexts. Gemini enables immediate deployment at \$30 annually (120K images $\times$ \$0.00025 within free-tier), achieving estimated 50–70% accuracy on common species with substantial online presence (lions, elephants, common birds) but potentially 30–40% accuracy on rare or endangered species with minimal photography (fewer than 100 online images). Despite reduced accuracy on rare species compared to our 67% performance on consumer products, this zero-shot capability proves valuable for exploratory research with manual verification workflows, representing the only economically viable approach given budget constraints. The architectural advantage proves decisive: many rare wildlife species completely lack representation in COCO training data, with YOLO fundamentally incapable of detecting these animals regardless of their rarity, while Gemini can leverage even limited web-crawled conservation imagery and scientific publications providing partial exposure during pre-training.
@@
Enterprise retail chains operating automated shelf monitoring across 1,000 stores with 500,000 images daily face dramatically different economics. Annual volume of 182.5 million images far exceeds 55.3 million break-even, enabling substantial annotation amortization. YOLO five-year TCO reaches \$75,492 (\$21,600 annotation + \$500 training + \$50,000 distributed edge hardware + \$3,392 maintenance) versus Gemini requiring \$228,125 (182.5M $\times$ 5 years $\times$ \$0.00025), yielding 67% cost savings for supervised approaches. Additionally, YOLO delivers 89.7% accuracy meeting stringent 90% requirements for inventory regulatory reporting, 9.2ms latency enabling real-time monitoring, and edge deployment eliminating network dependencies. Gemini’s 72.8% accuracy falls below compliance thresholds and 287ms latency prevents real-time operation, making supervised training both economically and technically superior.
Autonomous vehicle systems processing 10 million images daily prioritize supervised approaches despite annotation costs. Safety requirements mandate 95%+ accuracy achievable only through supervised training with extensive validation datasets, while sub-20ms latency requirements eliminate 287ms VLM approaches regardless of cost considerations. The \$140,500 one-time YOLO investment (20 categories $\times$ 5,000 images $\times$ 3 boxes $\times$ \$0.30 $\times$ 1.20 annotations + \$5,000 training + \$30,000 high-performance edge hardware + \$3,500 annual redundancy) proves negligible compared to liability exposure from detection failures, with per-vehicle costs of \$140 over five-year deployment lifetime acceptable for safety-critical systems. Gemini’s architectural unsuitability combining insufficient accuracy, excessive latency, and API dependency creating single point of failure eliminates consideration regardless of operational cost advantages.

### v2

@@ 5.1 Novel Category Evaluation with Tiered Web Representation
Novel category evaluation reveals fundamental architectural differences with critical insights from stratified web prevalence sampling. YOLOv8m detects exactly 0 of 500 novel product instances across all three tiers as architecturally expected, validating that supervised detector cannot generalize beyond training taxonomies regardless of web representation density, model quality, or training data scale.
@@
In stark contrast, Gemini successfully detects 262 of 500 instances (52.3% overall accuracy, CI [48.9%, 55.7%]) and GPT-4 detects 276 of 500 instances (55.1% overall accuracy, CI [51.7%, 58.5%]) with dramatic performance variation across web prevalence tiers:
@@
Tier 1 (Highly Web-Prevalent, N=250): Gemini 79.0% [75.2%, 82.6%], GPT-4 82.4% [78.9%, 85.6%] This tier includes consumer electronics with millions of online images (AirPods, Tesla, smart home devices). High performance reflects substantial VLM pre-training exposure through product marketing, user-generated content, e-commerce listings, and review platforms. Figure 7 shows some successful zero-shot detection via VLMs.
Per-category breakdown reveals nuanced performance:
@@
  • AirPods Pro: 90% (9/10) - Distinctive charging case design and ubiquitous Apple marketing enable reliable detection
@@
  • Ring Doorbell: 90% (9/10) - Prominent camera module and extensive installation video coverage
@@
  • Tesla Model 3: 80% (8/10) - Minimalist automotive design and extensive media coverage
@@
  • Amazon Echo Dot: 80% (8/10) - Cylindrical form factor and tech review prevalence
@@
  • Peloton Bike: 80% (8/10) - Distinctive screen mount and fitness media coverage
Figure 7: Successful zero-shot detection on web-prevalent products.
In this tier GPT-4 demonstrates a consistent 2-4 percentage point advantage over Gemini across most categories in this tier, suggesting superior capability for recognizing highly web-prevalent objects.
@@
Tier 2 (Moderately Web-Prevalent, N=150): Gemini 48.0% [42.8%, 53.2%], GPT-4 51.3% [46.0%, 56.6%] This tier includes products with thousands rather than millions of online images. Moderate performance reflects partial VLM exposure:
@@
  • Electric Scooters: 60% (6/10) - Vehicle-like characteristics and urban mobility coverage
  • N95 Masks: 50% (5/10) - Medical equipment similarity and pandemic surge imagery
  • Standing Desks: 50% (5/10) - Furniture similarities and office equipment niche
@@
  • Smart Door Locks: 40% (4/10) - Hardware similarities and home security niche
@@
  • Portable Power Stations: 40% (4/10) - Battery pack similarities and outdoor niche
Figure 8: Variable performance on moderately prevalent products.
As shown in Figure 8, GPT-4 maintains 3-4 percentage point advantage over Gemini in this tier, with both models struggling on categories with minimal distinctive visual features.
@@
Tier 3 (Rare Specialized Equipment, N=100): Gemini 30.0% [24.2%, 35.8%], GPT-4 32.0% [26.0%, 38.0%] This tier includes equipment with fewer than 500 online images, testing genuine zero-shot capability. Figure 9 shows a successful detection example for this tier. Low performance confirms limited pre-training exposure:
@@
  • Industrial 3D Metal Printers: 20% (2/10) - Specialized manufacturing equipment with confidential documentation
@@
  • Underwater ROV Equipment: 30% (3/10) - Marine research niche with limited public imagery
@@
  • Precision Laboratory Pipettes: 40% (4/10) - Scientific equipment with rare lab photography
  • Professional Broadcast Cameras: 30% (3/10) - Specialized models in production environments
@@
  • Medical Ultrasound Transducers: 30% (3/10) - Clinical equipment with privacy restrictions
Figure 9: True zero-shot capability on rare specialized equipment.
GPT-4 shows marginal 2 percentage point advantage over Gemini, suggesting both models struggle with truly rare objects, though this performance substantially exceeds YOLO’s architectural impossibility (0%).
@@
These tiered results validate our hypothesis: VLM zero-shot performance directly correlates with web representation density, ranging from near-supervised levels (79-82%) on web-prevalent products to genuine zero-shot challenges (30-32%) on rare equipment. However, even 30-32% substantially exceeds YOLO’s architectural impossibility (0%), establishing decisive qualitative advantage for applications encountering any objects outside training distributions.
The statistical reliability of these measurements is reinforced by high inter-annotator agreement (Fleiss’ kappa: 0.91 Tier 1, 0.87 Tier 2, 0.81 Tier 3) and comprehensive sample coverage across 50 diverse categories.
Figure 10 visualizes the tiered performance summary, illustrating the detection propensity for each of the VLMs at the three levels.
Figure 10: Zero-shot performance correlation with web representation density.
@@ 6 Cost-Effectiveness Analysis and Deployment Scenarios
Integration of technical performance results with cost modeling framework enables comprehensive economic analysis revealing when each paradigm achieves optimal return on investment. Table 3 presents TCO projections and Cost per correct detection calculations at representative inference volumes for 100-category systems, demonstrating how cost-effectiveness evolves with deployment scale.Figure 11 shows the CCD evolution for all the models that were used.
Table 3: Comparative inference cost and efficiency between YOLO, Gemini, and GPT-4
Inference YOLO YOLO Gemini Gemini GPT-4 GPT-4 Best
Volume TCO (\$) CCD (\$) TCO (\$) CCD (\$) TCO (\$) CCD (\$) Choice
1K 11,620 12.96 0.25 0.00034 10.00 0.00135 Gemini
10K 11,624 1.30 2.50 0.00034 100.00 0.00135 Gemini
100K 11,658 0.130 25.0 0.00034 1,000.0 0.00135 Gemini
1M 12,072 0.0135 250 0.00034 10,000 0.00135 Gemini
10M 15,752 0.0018 2,500 0.00034 100,000 0.00135 Gemini
50M 30,032 0.0007 12,500 0.00034 500,000 0.00135 Gemini
100M 47,632 0.00053 25,000 0.00034 1,000,000 0.00135 Gemini
150M 65,232 0.00048 37,500 0.00052 1,500,000 0.00192 YOLO
200M 82,832 0.00046 50,000 0.00069 2,000,000 0.00256 YOLO
Figure 11: Cost per correct detection of evolution across deployment scales.
@@
Despite Gemini’s 22.7 percentage point and GPT-4’s 19.9 percentage point accuracy disadvantages, the per-detection cost advantage remains substantial at this scale.
As deployment volume increases toward break-even thresholds, annotation costs amortize across growing detection counts, improving supervised cost-effectiveness. At 10 million inferences, YOLO CCD improves to \$0.0018 but still exceeds Gemini’s constant \$0.00034 and GPT-4’s constant \$0.00135 by 5.3-fold and 1.3-fold respectively.
@@
Only beyond approximately 120–150 million inferences does YOLO achieve decisively superior CCD when accounting for 89.7% accuracy, delivering 17.9 million more successful detections than Gemini’s 72.8% accuracy at that scale.
These break-even volumes far exceed typical deployment scales for most organizations. A startup processing 1,000 images daily reaches only 365,000 annual inferences, operating at 0.7% of TCO break-even and 0.3% of CCD break-even. A small-to-medium business processing 5,000 images daily reaches 1.825 million annually (3.3% of break-even).
Even substantial enterprises processing 50,000 images daily require 3.0 years continuous operation to approach 55 million inference TCO break-even, with realistic deployments facing evolving requirements and hardware refreshes preventing indefinite amortization. The practical implication establishes that supervised training proves economically rational only for the highest-volume production systems processing hundreds of thousands to millions of images daily across stable category taxonomies.
@@
Scenario Daily Cat. Budget Acc. Lat. Optimal Primary Rationale
Vol
@@
Startup E-commerce 1K 50 <<\$5K 65%+ Any Gemini YOLO \$6K upfront exceeds budget; Gemini \$0 upfront + \$90/year
@@
SMB Retail Analytics 5K 100 \$10K 75%+ <<1s Gemini Annual vol 1.8M far below 55M break-even; Gemini \$456/year vs YOLO \$11K upfront
@@
Research Wildlife 333 500+ <<\$3K 60%+ Any GPT-4 YOLO \$180K annotation impossible; novel species require zero-shot
@@
Medical Imaging 10K 12 \$50K 85%+ <<10s Hybrid Expert annotation \$43K prohibitive; hybrid VLM screening + YOLO verification
@@
Enterprise Inventory 500K 200 \$500K 90%+ <<50ms YOLO Annual vol 182.5M exceeds break-even; YOLO 5-yr \$75K vs Gemini \$228K
@@
Autonomous Vehicles 10M 20 Unlim. 95%+ <<20ms YOLO Safety demands 95%+ accuracy; 9.2ms latency mandatory
@@
An e-commerce startup deploying detection across 50 product categories initially, with planned expansion to additional categories, processing approximately 1,000 images daily faces YOLO upfront costs of \$6,020 (annotation \$5,400 + training \$120 + hardware \$500) versus zero for Gemini and \$10 for GPT-4. Note that most startups begin with 5–10 core categories and expand incrementally; our 50-category scenario represents a more aggressive initial deployment.
With annual operational expenses of \$6,020 for supervised versus \$90 for zero-shot (365K images $\times$ \$0.00025 minus free-tier coverage) and \$1,314 for GPT-4 (365K images $\times$ \$0.01 minus free-tier coverage), the supervised approach exceeds typical startup budgets under \$5,000 while requiring 3–4 weeks for annotation completion, making Gemini the economically viable option for rapid market entry.
Gemini delivers 68.5% accuracy and GPT-4 delivers 71.3% accuracy exceeding typical 65% threshold for catalog applications where human review provides backup verification. Additionally, category evolution at 5–10 new products monthly imposes \$144 per category (\$90 annotation + \$54 training) for YOLO versus \$0 for VLMs, further strengthening zero-shot advantages for rapidly evolving taxonomies common in e-commerce.
@@
Small-to-medium businesses operating retail analytics across multiple stores with 5,000 images daily face annual volume of 1.825 million images, reaching only 3.3% of the 55.3 million break-even thresholds. YOLO requires \$11,616 upfront investment for 100-category system versus Gemini annual cost of \$456 (after free-tier deduction) and GPT-4 annual cost of \$18,250, yielding 25-fold cost advantage for Gemini over YOLO and 40-fold advantage over GPT-4.
While YOLO provides superior 91.2% accuracy versus Gemini’s 68.5% and GPT-4’s 71.3%, the accuracy improvements cost \$11,160 additional over first year for Gemini and \$6,366 for GPT-4, equivalent to \$657 and \$375 per percentage point of accuracy gain respectively. For retail analytics applications where 75% accuracy suffices with occasional manual verification, this premium proves economically unjustifiable.
Research laboratories exploring wildlife monitoring face annotation costs scaling with species count. For a realistic initial study targeting 50 species with 10,000 monthly images (333 daily average), annotation costs reach \$18,000 (50 categories $\times$ 100 images $\times$ 3 boxes $\times$ \$0.30 $\times$ 1.20), already exceeding typical grant budgets of \$5,000–10,000 for pilot projects. Expanding to comprehensive biodiversity monitoring across 500+ species would require \$180,000 annotation investment, prohibitive for most research contexts.
Gemini enables immediate deployment at \$30 annually (120K images $\times$ \$0.00025 within free-tier), while GPT-4 would cost \$1,200 annually. Given the higher accuracy of GPT-4 (55.1% vs 52.3% on novel categories) and the importance of accuracy in research contexts, GPT-4 represents the optimal choice for research applications despite higher costs.
Both VLMs achieve estimated 50–70% accuracy on common species with substantial online presence (lions, elephants, common birds) but potentially 30–40% accuracy on rare or endangered species with minimal photography (fewer than 100 online images). Despite reduced accuracy on rare species compared to our performance on consumer products, this zero-shot capability proves valuable for exploratory research with manual verification workflows, representing the only economically viable approach given budget constraints.
The architectural advantage proves decisive: many rare wildlife species completely lack representation in COCO training data, with YOLO fundamentally incapable of detecting these animals regardless of their rarity, while VLMs can leverage even limited web-crawled conservation imagery and scientific publications providing partial exposure during pre-training.
@@
Enterprise retail chains operating automated shelf monitoring across 1,000 stores with 500,000 images daily face dramatically different economics. Annual volume of 182.5 million images far exceeds 55.3 million break-even, enabling substantial annotation amortization. YOLO five-year TCO reaches \$75,492 (\$21,600 annotation + \$500 training + \$50,000 distributed edge hardware + \$3,392 maintenance) versus Gemini requiring \$228,125 (182.5M $\times$ 5 years $\times$ \$0.00025) and GPT-4 requiring \$9,125,000 (182.5M $\times$ 5 years $\times$ \$0.01), yielding 67% cost savings for supervised approaches over Gemini and over 120-fold savings over GPT-4.
Additionally, YOLO delivers 91.2% accuracy meeting stringent 90% requirements for inventory regulatory reporting, 9.2ms latency enabling real-time monitoring, and edge deployment eliminating network dependencies. Gemini’s 68.5% accuracy and GPT-4’s 71.3% accuracy fall below compliance thresholds and their latencies prevent real-time operation, making supervised training both economically and technically superior.
Autonomous vehicle systems processing 10 million images daily prioritize supervised approaches despite annotation costs. Safety requirements mandate 95%+ accuracy achievable only through supervised training with extensive validation datasets, while sub-20ms latency requirements eliminate 287ms Gemini and 312ms GPT-4 approaches regardless of cost considerations.
The \$140,500 one-time YOLO investment (20 categories $\times$ 5,000 images $\times$ 3 boxes $\times$ \$0.30 $\times$ 1.20 annotations + \$5,000 training + \$30,000 high-performance edge hardware + \$3,500 annual redundancy) proves negligible compared to liability exposure from detection failures, with per-vehicle costs of \$140 over five-year deployment lifetime acceptable for safety-critical systems.
VLMs’ architectural unsuitability combining insufficient accuracy, excessive latency, and API dependency creating single point of failure eliminates consideration regardless of operational cost advantages.